OpenStudy (anonymous):

assume that a very long rope supported by poles that are 3 ft tall is stretched around the moon. neglect gravity pull and rope takes the shape of a circle). The radius of the moon is approx. 1080miles. How much longer would you need to make the rope if you wanted the rope supported by poles that are 4 ft tall. (use 3.14 for pi)

OpenStudy (anonymous):

ME TOO!

OpenStudy (anonymous):

rope wraps around them oon like a circle = circumference..? and How is the answer 6.28 ft lol

OpenStudy (anonymous):

jinx use I tried the formula and got a big retricenumber

OpenStudy (anonymous):

please someone if i dont get this problem My teacher will burn all of my section homework

OpenStudy (anonymous):

c =2(3.14)r is circum

OpenStudy (aum):

Circumference = \(2\pi r\) The difference in circumference = \(2\pi r_2 - 2\pi r_1 = 2\pi(r_2-r_1) = \\ 2\pi(4-3) =2\pi(1) = 2 * 3.14 = 6.28 \text{ feet.} \)

OpenStudy (aum):

When we take the difference in radius, the radius of the moon will cancel out leaving with just the difference in the height of the poles which is 1 foot.